OpenStudy (vortish):

find the verical asymptotes if any and the values if any of (x) and the corresponding holes if any of the rational graph f(x)= x^2-16/x-4

OpenStudy (freckles):

well we have a fraction the fraction is undefined when the bottom is 0 when is the bottom 0?

OpenStudy (vortish):

because x-4 is x+4

OpenStudy (vortish):

which = 0

OpenStudy (freckles):

x-4 isn't x+4 i'm asking when the bottom is 0 when is x-4=0

OpenStudy (vortish):

then I dont know

OpenStudy (freckles):

you never solve equations like x-4=0 before?

OpenStudy (freckles):

i will give you a hint add 4 on both sides

OpenStudy (freckles):

x-4+4=0+4

OpenStudy (freckles):

what is -4+4?

OpenStudy (vortish):

0

OpenStudy (freckles):

so x+0=0+4 so x=4 x-4=0 when x=4 the bottom is 0 when x=4

OpenStudy (freckles):

that means there is a hole at x=4 or a vertical asymptote

OpenStudy (freckles):

to figure out which one we need to factor the top and see if we can get the bottom factor to cancel out

OpenStudy (freckles):

can you factor x^2-16?

OpenStudy (vortish):

ok but what about the x^2-16

OpenStudy (vortish):

i know that 4 goes in to 16 4 times

OpenStudy (vortish):

but the x^2 has me stumpt

OpenStudy (freckles):

x^2-16 is a difference of squares to factor a^2-b^2 we get (a-b)(a+b)

OpenStudy (vortish):

ok

OpenStudy (vortish):

i hate these i never end up understanding a dam thing

OpenStudy (freckles):

well right now i'm just asking you to factor x^2-16

OpenStudy (freckles):

you can use the formula i gave if you want

OpenStudy (vortish):

that formula is were i always can no figure things out

OpenStudy (vortish):

x=0 (x-16)(x+16) ?

OpenStudy (freckles):

x^2-16=x^2-4^2=(x-4)(x+4)

OpenStudy (freckles):

\[f(x)=\frac{x^2-16}{x-4}=\frac{(x-4)(x+4)}{x-4}\] if we are able to cancel the x-4 on bottom then we have a hole at x=4 but if we we can't cancel the x-4 on bottom then we have a vertical asymptote at x=4 but as you notice it does cancel